# Scaling Efficient Masked Image Modeling on Large Remote Sensing Dataset

### v4

@@ 4 Experiment
Table 1: Performance comparison of using the MillionAID and OpticalRS-13M datasets separately with the same data throughput, evaluated by MAE [28] and ViT-B [19].
Dataset Method Image Number Epoch Scene Object Semantic
Classification Detection Segmentation
AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MillionAID [48] MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99 75.20 66.99 52.56 79.43
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60 74.50 66.36 52.63 79.39
@@
In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments. We also conducted further experiments to test the scalability of the pipeline in the ablation experiments. Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks. All experiments are performed by fine-tuning pre-trained models on downstream tasks, including RS scene classification, object detection, and semantic segmentation. The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
Table 2: The performance comparison between SelectiveMAE and the baseline method MAE [28] with the ViT-B backbone. The training time is measured on NVIDIA A100 GPUs. To accelerate validation, we select 4 million images from OpticalRS-13M for pre-training.
Method Image Number Epoch Total Training Scene Object Semantic
Classification Detection Segmentation
Time (h) AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
MAE 4 million 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE 4 million 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
@@ 4.1 Ablation Study
To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated training method, SelectiveMAE. Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes. More ablation experiment results related to model design are presented in the supplementary material.
Dataset Diversity of OpticalRS-13M. In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance. Here, We define dataset diversity as the ability to improve performance across various types of tasks. While datasets like MillionAID [48] focus on classification, our OpticalRS-13M dataset is designed to extend the benefit to detection and segmentation tasks. We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput. As shown in Table 1, even with random sampling of 1 million images from OpticalRS-13M, it outperforms MillionAID on downstream tasks, particularly in object detection and segmentation. The results in Table 1 also indicate that the optimal outcomes at equivalent data throughput are achieved by different settings other than using all 13 million images. This suggests that training on a large number of samples with few epochs was insufficient for the MIM method to fully leverage the large dataset, leading to inferior generalization on downstream tasks. Experimental results in the supplementary material show consistent performance improvements across downstream tasks with more training epochs.
@@
Table 3: The efficiency comparison between SelectiveMAE and the baseline method MAE [28] across different backbones on the MillionAID dataset [48] with 1600 epochs. The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
Model Backbone Params Total Training Time (h) GPU Memory (MB)
MAE ViT-B 86M 101 17628
SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×) 9570 (1.8×1.8$\times$1.8 ×)
MAE ViT-L 307M 113 30530
SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×) 18790 (1.6×1.6$\times$1.6 ×)
Performance Advantage of SelectiveMAE. To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-13M using both MAE [28] and SelectiveMAE. The performance of ViT-B on various downstream tasks after pre-training is presented in Table 2. The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed. This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
@@
Efficiency Advantage of SelectiveMAE. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pretraining of different backbones on the MillionAID dataset, as shown in Table 3 (the corresponding accuracies are presented in Table 5). It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models. Additionally, it is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches.
Scalability of the Pipeline To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each. The results in Table 4 show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours. When the dataset was expanded to 4 million samples, the time savings increased to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
Table 4: The training cost of SelectiveMAE and the baseline method MAE [28] on OpticalRS-13M with different data volumes. All experiments are conducted on NVIDIA A100 GPUs.
Method Images Number Epoch Total Training Time (h)
MAE 1 million 800 54
SelectiveMAE 1 million 800 25 (-29)
MAE 4 million 800 177
SelectiveMAE 4 million 800 86 (-81)
Table 5: Performance comparison results of different models. “TR” represents the ratio of training data to the entire dataset. The first and second scores are marked by bold and blue, respectively. * means pre-training on 4 million images sampled from OpticalRS-13M.
Model Backbone Params (M) Scene Object Semantic
Classification Detection Segmentation
AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [49] ResNet-50 26 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [1] ResNet-50 26 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [65] ResNet-50 26 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [50] ResNet-50 26 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [17] ViT-L 307 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [56] ViT-L 307 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [79] ViT-S 22 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [63] Swin-B 88 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [5] Swin-B 88 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [51] Swin-B 88 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [70] ViT-B+RVSA 86 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
LSKNet [44] LSKNet-S 14 97.05/98.22 94.27/95.83 - - 54.00 -
OREOLE [18] ViT-G 914 96.71/   -    -   /   - 77.40 71.31 54.00 -
SelectiveMAE* ViT-B 86 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE* ViT-L 307 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
SelectiveMAE ViT-B 86 96.64/98.30 92.53/94.50 75.10 67.69 52.68 79.44
SelectiveMAE ViT-L 307 97.21/98.48 94.16/95.90 78.70 71.75 53.92 79.48
4.2 Main Results
We compare SelectiveMAE to state-of-the-art RSFMs. In addition to benchmarking against MIM-based approaches such as SatMAE [17] and ScaleMAE [56], we also compare with contrastive learning-based methods, including GASSL [1] and SeCo [49], as well as advanced supervised learning models like SatLas [5]. Notably, none of these comparative methods were specifically designed for accelerated pre-training. Although their speed performance was not reported, we expect them to lag behind the baseline MAE.
Scene Classification. We first assess the pre-trained model’s performance on the scene classification task to evaluate its overall representational capability without additional decoders. We use two scene classification datasets: AID [86] and RESISC-45 [13], following training details and train-test splits as in [70, 56]. Table 5 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline. Additionally, it surpasses LSKNet and others when scaled to ViT-L, indicating that, with SelectiveMAE on the OpticalRS-13M dataset, the model effectively learns strong feature representations and scales efficiently.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [41] and its improved variant DIOR-R for oriented object detection [15]. Following the methodologies of previous works [70, 63], we maintained consistent experimental setups, employing Faster-RCNN [57] and Oriented-RCNN [88] as detectors for each dataset. The results are summarized in Table 5. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [63]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE, which has close to 1B parameters, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [76] and SpaceNetv1 [68]. Our implementation follows [70], utilizing UperNet [87] as the segmentation framework. Table 5 demonstrates the clear superiority of SelectiveMAE over its competitors in semantic segmentation tasks. SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
@@ 5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. Built upon it, we first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-13M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that using only 40% RS image patches is sufficient for training a comparable MIM model. Extensive experiments show that OpticalRS-13M significantly contributes to improving classification, detection, and segmentation performance, while SelectiveMAE achieves over a 2×2$\times$2 × speedup along with GPU memory savings, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.

### v5

@@ 4 Experiment
In this section, we performed comprehensive comparative experiments to evaluate the performance of the proposed method on various downstream tasks, including RS scene classification, object detection, and semantic segmentation. Then, we validated the effectiveness of the OpticalRS-13M dataset and the SelectiveMAE method through ablation experiments. We also conducted further experiments to test the scalability of the pipeline. All experiments are performed by fine-tuning. The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
Table 1: Performance comparison results of different models. “TR” represents the ratio of training data to the entire dataset. The first and second scores are marked by bold and blue, respectively. $\dagger$ means pre-training on 4 million images sampled from OpticalRS-13, with the epoch is set to 800. The overall accuracy is adoped as the metric for the scene classification task.
Model Backbone Params (M) Data Scene Object Semantic
Throughput Classification Detection Segmentation
/Minute AID [93] RESISC-45 [94] DIOR [1] DIOR-R [95] LoveDA [55] SpaceNetv1 [56]
TR=20%/50% TR=10%/20% mAP50 mAP50 mIoU mF1
SeCo [25] ResNet-50 [96] 26 - 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [27] ResNet-50 [96] 26 - 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [31] ResNet-50 [96] 26 - 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [24] ResNet-50 [96] 26 - 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [33] ViT-L [2] 307 205k 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [37] ViT-L [2] 307 206k 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [58] ViT-S [2] 22 - 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [41] Swin-B [97] 88 - 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [60] Swin-B [97] 88 243k 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [28] Swin-B [97] 88 - 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [21] ViT-B+RVSA [21] 86 - 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
OREOLE [43] ViT-G [43] 914 - 96.71/   -    -   /   - 77.40 71.31 54.00 -
MAE [13$\dagger$ ViT-B [2] 86 264k 96.58/98.02 92.44/94.43 75.40 67.35 52.80 79.41
SelectiveMAE $\dagger$ ViT-B [2] 86 556k 96.90/98.12 93.35/94.58 75.70 67.78 53.05 79.50
[…]
SelectiveMAE ViT-B [2] 86 556k 97.10/98.28 93.70/95.48 75.80 67.69 52.68 79.44
SelectiveMAE ViT-L [2] 307 533k 97.49/98.52 94.73/96.36 78.70 71.75 53.92 79.48
4.1 Main Results
We compare SelectiveMAE to state-of-the-art RSFMs. In addition to benchmarking against MIM-based approaches such as SatMAE [33] and ScaleMAE [37], we also compare with contrastive learning-based methods, including GASSL [27] and SeCo [25], as well as advanced supervised learning models like SatLas [60]. Notably, these comparative methods were specifically designed for increasing performance rather than accelerating pre-training. The Scene Classification and Detection comparison results are taken directly from SkySense [83], while those for Semantic Segmentation are exactly sourced from MTP [38]. Although their speed performance was not reported, we expect them to lag behind the baseline MAE. The performance of ViT-B and ViT-L on various downstream tasks after pre-training is presented in Table 1. The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE outperforms other RSFMs. Notably, it not only outperforms baseline (MAE) but also achieves this at 2.1 times the speed (556/264). This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
Scene Classification. We first assess the pre-trained model’s performance on the scene classification task to evaluate its overall representational capability without additional decoders. We use two scene classification datasets: AID [93] and RESISC-45 [94], following training details and train-test splits as in [21, 37]. Table 1 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets. Additionally, when scaled to ViT-L, our model outperforms OREOLE [43] and other competitors, indicating the efficacy of SelectiveMAE pre-training on the OpticalRS-13M dataset, which enables strong feature representation learning while maintaining efficient scalability with increasing model size.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [1] and its improved variant DIOR-R for oriented object detection [95]. Following the methodologies of previous works [21, 41], we maintained consistent experimental setups, employing Faster-RCNN [98] and Oriented-RCNN [99] as detectors for each dataset. The results are summarized in Table 1. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [41]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE [43], which has close to 1B parameters, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known RS datasets: LoveDA [55] and SpaceNetv1 [56]. Our implementation follows [21], utilizing UperNet [100] as the segmentation framework. Table 1 demonstrates the clear superiority of SelectiveMAE over its competitors in semantic segmentation tasks. SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
@@ 4.2 Ablation Study
To validate the efficacy of SelectiveMAE, we conducted ablation experiments about HOG Selecting, Similarity Measure and Selection Pattern. To further verify the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated performance of SelectiveMAE. Additionally, we compared the SelectiveMAE with the CrossMAE, a MAE-based method for natural images. More ablation experiment results are presented in the supplementary material.
Alternatives of HOG Selecting. To verify the efficacy of HOG in token selecting, we conduct two ablation experiments using MillionAID dataset [23] to replace HOG: (1) we used a pretrained Swin-B [97] model and applied k-nearest neighbor clustering to the extracted features. (2) We employed a lightweight feature extractor network based on Adamae [101]. Results in the Table 2 demonstrate HOG’s pre-training speed advantage. While using Swin-B slightly outperforms HOG, it is highly inefficient, and pre-training a foundation model like Swin requires massive data and computational resources. For RS images with abundant low-level features, HOG effectively selects tokens while ensuring substantially higher training speeds.
Table 2: Ablation experiment results of HOG selection through pre-training on the MillionAID dataset for 800 epochs.
Method Params Data Throughput AID RESISC-45
/ Minute TR=20%/50% TR=10%/20%
Adamae [101] 2.36M 498k 88.78/91.25 85.72/87.44
Swin-B 88M 356k 93.21/96.48 89.94/93.72
HOG - 556k 93.17/96.12 89.21/92.31
Ablation of Similarity Measure and Selection Pattern. To comprehensively assess the PSTS, we tested different Similarity Measures and the Selection Pattern. The results in the Table 3 reveal that different similarity measures had minimal impact on performance. However, changing the selection pattern led to a performance drop, confirming the effectiveness of the PSTS.
Table 3: Ablation experiment results of different similarity measures and selection patterns in PSTS through pre-training on the MillionAID dataset for 800 epochs.
Selection Pattern Similarity Measure AID RESISC-45
TR=20%/50% TR=10%/20%
near-far-random Euclidean distance 93.02/95.96 88.97/92.07
near-far-random Manhattan distance 92.92/95.89 89.17/91.92
far-near-random Cosine distance 90.12/92.78 85.81/88.80
near-far-random Cosine distance 93.17/96.12 89.21/92.31
Table 4: Performance comparison of using different pre-training settings on datasets, methods and epochs, where the ViT-B [2] is adopted as the network.
Dataset Method Image Number Epoch AID [93] RESISC-45 [94]
TR=20%/50% TR=10%/20%
Different Datasets
MillionAID [23] MAE 1 million 800 94.92/97.38 89.20/93.60
OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88
Equivalent Data Throughput
OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31
OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41
OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30
OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60
OpticalRS-13M SelectiveMAE 1 million 800 96.29/97.78 91.41/93.48
OpticalRS-13M SelectiveMAE 4 million 200 95.96/98.06 92.06/94.05
OpticalRS-13M SelectiveMAE 13 million 67 96.31/97.95 91.88/93.76
More Epoch
OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60
OpticalRS-13M MAE 13 million 100 96.34/98.13 91.79/93.85
OpticalRS-13M MAE 13 million 200 96.51/98.22 92.19/94.26
OpticalRS-13M MAE 13 million 800 97.10/98.28 93.70/95.48
Efficacy of OpticalRS-13M.
As illustrated in Table 4, when pre-training with a randomly sampled subset of 1 million images from OpticalRS-13M, the model achieves superior performance compared to MillionAID [23], underscoring the efficacy of OpticalRS-13M for representation learning. To further assess dataset diversity, we conducted experiments with equivalent data throughput during pre-training. The results in Table 4 reveal that optimal performance is attained by varying training configurations rather than utilizing the entire 13 million image corpus, regardless of the pre-training methodology. In our opinion, this observation highlights the inherent diversity of OpticalRS-13M, as training with fewer epochs proved insufficient for the MIM method to fully exploit the dataset’s potential, leading to model underfitting. To validate this hypothesis, we extended the training schedule, and the experimental results presented in the last part of Table 4 confirm our claim.
@@
Table 5: The efficiency comparison between SelectiveMAE and the baseline method MAE [13] across different backbones on the part of OpticalRS-13M dataset [23] with 800 epochs. The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
Model Backbone Data Volume Training Time (h) GPU Memory (MB)
MAE ViT-B 1 million 51 17628
SelectiveMAE ViT-B 1 million 24 (2.1×2.1$\times$2.1 ×, -27) 9570 (1.8×1.8$\times$1.8 ×)
MAE ViT-L 1 million 57 30530
SelectiveMAE ViT-L 1 million 25 (2.3×2.3$\times$2.3 ×, -32) 18790 (1.6×1.6$\times$1.6 ×)
MAE ViT-B 4 million 177 17628
SelectiveMAE ViT-B 4 million 86 (2.1×2.1$\times$2.1 ×, -81) 9570 (1.8×1.8$\times$1.8 ×)
@@
Efficiency Advantage of SelectiveMAE. To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pre-training of different backbones on the part of OpticalRS-13M dataset, as shown in Table 5 (the corresponding accuracies are presented in Table 1). It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models. It is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches. To further evaluate the scalability of our pipeline, we organized datasets containing 4 million samples and conducted experiments. The results in Table 5 show that, when the dataset was expanded to 4 million samples, the time savings increased from 27 (51-24) to 81 hours. These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
Comparison with CrossMAE [49]. CrossMAE, an MAE-based self-supervised pretraining method, is originally designed for natural images. Since SelectiveMAE adopts the same decoder as CrossMAE [49] during pre-training, and incorporates a HOG-based strategy tailored to RS images inspired by its partial reconstruction concept, as detailed in Section 3.2.2. In this section, we have also complemented related comparison experiments. Nevertheless, due to CrossMAE has not been pretrained in the RS field, we conducted a fair evaluation through pretraining it on the OpticalRS-13M dataset using the official code. Table 6 shows that SelectiveMAE outperforms CrossMAE in both accuracy and efficiency. By integrating RS-specific enhancements, PSTS and partial reconstruction, SelectiveMAE proves to be a highly effective self-supervised learning approach for RS.
Table 6: Performance comparison between our method and CrossMAE, where 400 million images of OpticalRS-13M are randomly sampled for pre-training 800 epochs.
Method Data Throughput / Minute AID RESISC-45
TR=20%/50% TR=10%/20%
CrossMAE(ViT-B) 475k 96.12/97.18 92.08/94.12
SelectiveMAE(ViT-B) 556k 96.78/98.12 93.35/94.58
@@ 5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. We first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-13M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that using only 40% RS image patches is sufficient for training a comparable MIM model. Extensive experiments show that OpticalRS-13M significantly contributes to improving classification, detection, and segmentation performance, while SelectiveMAE achieves over a 2×2$\times$2 × speedup along with GPU memory savings, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.